# Symbolic Powers via Extension

### v2

@@ 2. Symbolic Powers via flat extension
Now by Lemma 2.4, we have $\bigcup_{\operatorname{Ass}{\mathfrak{q}}\in\left(I^{(n)}B\right)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(IB)}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B$. Thus whenever $s\in W$ then $s\not\in{\mathfrak{q}}$, where ${\mathfrak{q}}$ is an associated prime of $I^{(n)}B$. Thus by [AM94, Proposition 4.4(iii)] we get
@@
In Theorem 2.5 we extend [GHMN17, Proposition 3.8], [Ake17, Theorem 4], and [Wal18, Proposition 2.1] by replacing the faithfully flat condition with weaker conditions. Note that faithfully flatness will implies the assumptions in Theorem 2.5. But not vice versa.
@@
We give simple examples satisfying assumptions in Theorem 2.5 but not in [Wal18, proposition 2.1]. Let $A$ be a ring having at least two different prime ideals such that one is not contained in another one, and $B=A_{\mathfrak{p}}$ be localization of $A$ at an arbitrary prime $\mathfrak{p}$. It is a well-known fact that localization is an exact functor, so $B$ is a flat $A$-module, and prime ideal $\mathfrak{p}$ in $A$ is the contraction of $\mathfrak{p}B$. Note that $\mathfrak{p}B$ is the only prime in $B$, then $\operatorname{Spec}(B)\rightarrow\operatorname{Spec}(A)$ is never surjective, so $B$ is not faithfully flat as a $A$-module according to [AM94, ex 3.16].
@@
One can take $A={\mathbb{Z}}$ and $B$ as $\mathbb{Z}_{(p)}$ where $p$ is a prime number. Take $I=(p^{m})$ for some $m\geqslant 1$. Then $I$ satisfies assumptions in Theorem 2.5. Furthermore, letting $B$ be any localization of $A$ will meet the assumptions in Theorem 2.5. This tells us that the stalk of $\operatorname{Spec}A$ gives us enough information of symbolic power of an ideal generated by the corresponding point in $\operatorname{Spec}A$.
Another example would be take $A={\mathbb{C}}[t]$ and $B={\mathbb{C}}[t]_{(p(t))}$, where $p(t)$ is an irreducible polynomial in ${\mathbb{C}}[t]$.
3. Resurgence and Prime extension
Definition 3.1.
A ring homomorphism $\phi:A\to B$ has prime extension property if for every prime $\mathfrak{p}$ of $A$, $\phi({\mathfrak{p}})=\mathfrak{p}B$ is also prime in $B$. We also call $\phi$ is a prime extension.
Remark 3.2.
We recall some facts about the morphisms with prime extension properties.
  1. (1)
    In many literature, (for example, see, [HJ21]) prime extension has been defined as either ${\mathfrak{p}}B$ is a prime ideal or, the unit ideal. For us, prime extension means that prime ideals extends to a prime ideals.
  2. (2)
    Flat and prime extension implies faithfully flatness. Indeed, by [Mat89, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat.
Proposition 3.3.
@@ 3. Resurgence and Prime extension
Let $A$ and $B$ be Noetherian rings and $\phi:A\to B$ be a flat morphism. If $I\subset A$ is a proper ideal and $\phi$ has prime extension properties then $\rho(I)=\rho(IB)$ and $\rho_{a}(I)=\rho_{a}(IB)$.
@@
Since $\phi$ is prime extension then by remark 3.2, it is faithfully flat. Thus using Theorem 2.5 and [AM94, ex 3.16(I)] we get that
$I^{(m)}\subset I^{r}\text{ if and only if }(IB)^{(m)}\subset(IB)^{r},\text{ for all }m,r.$
Hence
$\rho(I)=\rho(IB),\text{ and }\rho_{a}(I)=\rho_{a}(IB).$
We want to mention that [Wal18, Proposition 2.1] also proved the similar result.
Flat and prime extension has geometric meanings. Before we start to talk about them, we recall the definition.
@@
$B_{\mathfrak{p}}\text{ is an integral domain }\iff B/\mathfrak{p}B\text{ is anintegral domain }\iff\mathfrak{p}B\text{ is prime}.$
This proposition implies that if $A$ is reduced, and $\phi$ has prime extension property, then $B$ is also reduced. Indeed, if $A$ contains the zero ideal as a prime ideal, so does $B$. It also gives us an enormous class of examples. If $\phi$ is a prime extension, then every $B_{\mathfrak{p}}$ is an integral domain, which is geometrically interesting. Before we give an example, we recall the following lemma.
@@
[Sta24, Lemma 10.39.5] Let $M$ be an $R$-module, then $M$ be a flat $R$-module if only if for every ideal $I\subset R$ the map $I\otimes_{R}M\rightarrow R\otimes_{R}M$ is injective.
@@
We identify closed points of $A$ with elements of ${\mathbbm{k}}$, and obviously the $(t-c)$ in ${\mathbbm{k}}[t,x,y]/(y-x^{2}-tx)$ is irreducible for every $c\in{\mathbbm{k}}$, so $\phi$ is also a prime extension. The fiber $\operatorname{Spec}B_{(t-c)}$ is given by the plane curve $y=x^{2}+cx$ in $\mathbb{A}_{\mathbbm{k}}^{2}$, which is an irreducible variety, and those plane curves are even homeomorphic to each other.
The situation doesn’t always perform so nicely. If we consider $A$ as ${\mathbbm{k}}[t,s]$ and $B$ as ${\mathbbm{k}}[t,s,x,y]/(y-x^{2}-tx)$, we also have $(t-c)$ as our point, then $B\otimes_{A}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ will be ${\mathbbm{k}}[s,s^{-1},x,y]/(y-x^{2}-cx)={\mathbbm{k}}(s)[x,y]/(y-x^{2}-cx)$. However we have $(t-a,s-b)$ as our point where $a$,$b\in{\mathbbm{k}}$, now the corresponding fiber is $\operatorname{Spec}{\mathbbm{k}}[x,y]/(y-x^{2}-ax)$, which is not homeomorphic to the previous one.
@@ 4. Resurgence of sum of ideals
Reviewing the recent research of resurgence number of sums of ideals for polynomial rings over fields, our goal is to prove these results for sums of finitely generated ${\mathbbm{k}}$-algebras which are integral domains where ${\mathbbm{k}}$ is algebraically closed. In [BHJT21, Theorem 2.6] it has been proved that if $I$ and $J$ are ideals in different polynomial rings then
@@
In [KNT24, Theorem 2.3], it has been proved that
@@
In this section we prove similar results for finitely generated regular ${\mathbbm{k}}$-algebra domains, where ${\mathbbm{k}}$ is an algebraically closed field. One of the key tools in [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3] is the binomial expansion of symbolic powers of ideals, which holds for ideals in ${\mathbbm{k}}$-algebras. The other tool is $\rho(IR)=\rho(I)$, this holds in the polynomial ring set up. To generalize [BHJT21, Theorem 2.6] and [KNT24, Theorem 2.3], we need that $\rho(IR)=\rho(I)$ in general domains, which is not known.
In Lemma 4.1, we prove that if $A$ and $B$ are finitely generated algebra domains over an algebraically closed field ${\mathbbm{k}}$, and $R=A\otimes B$, then $\rho(IR)=\rho(I)$, and $\rho(JR)=\rho(J)$, where $I\subset A,J\subset B$ are ideals. In this section, we assume that all things are regular.
@@
Let $A$ and $B$ be finitely generated ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed, which are integral domains. If $\phi:A\to R$ is a canonical ring homomorphism, where $R=A\otimes B$, then $\phi$ is flat and has prime extension property. Furthermore, $\rho(IR)=\rho(I)$ and $\rho_{a}(IR)=\rho_{a}(I)$.
@@
Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [Iit82, Lemma 1.54]. It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property. The last equality follows from Proposition 3.3. ∎
Remark 4.2.
To see why we need the conditions on $A$ and $B$, we will need some examples and theorems. Now [Vám78, Theorem 11] implies that, if $A$ (as a field extension over ${\mathbbm{k}}$) is not finitely generated as a ${\mathbbm{k}}$-algebra, then $A\otimes_{\mathbbm{k}}A$ is not a Noetherian ring. For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring. Therefore, the condition being finitely generated is necessary. We also need $A$ and $B$ to be integral domains as ${\mathbbm{k}}$-algebras where ${\mathbbm{k}}$ is algebraically closed because of [Iit82, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain. It’s a famous example that $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$ is not an integral domain. Also note by [TY03, Theorem 6(e)] $A\otimes B$ is regular.
The next lemma also extends [BHJT21, Lemma 2.4]. The proof of [BHJT21, Lemma 2.4] depends on divisibility in the polynomial ring. While working with general domains, we treated them as vector spaces and used properties of basis.
@@
Take $\theta\in{\mathbb{Q}}$ such that $\theta>\rho_{a}^{\lim\sup}(I+J)$. Then there is a $s_{0}\in{\mathbb{N}}$ such that for all $t\geqslant s_{0},\theta>\rho(I+J,t)$, where $\rho(I+J,t)=\sup\left\{\dfrac{m}{r}~{}|~{}I^{(m)}\not\subset I^{r},m\geqslant t,r\geqslant t\right\}$. Thus for any $h,r\in{\mathbb{N}}$ so that $h,r\geqslant t\geqslant s_{0}$ and $\dfrac{h}{r}\geqslant\theta$, one has $(I+J)^{r}\supset(I+J)^{(h)}\supset I^{(h)}$. Up to this point the proof idea is same as [BHJT21, Lemma 2.4]. Now we need to show that $I^{(h)}\subseteq I^{r}$.
Let $V$ be a ${\mathbbm{k}}$ basis of $I$, which extends to $V^{*}$ a ${\mathbbm{k}}$-basis of $A$ containing $1$. Let $W$ be a ${\mathbbm{k}}$-basis of $J$, which extends to $W^{*}$, a ${\mathbbm{k}}$-basis of $B$ containing $1$. Then $V\otimes W^{*}$ is a ${\mathbbm{k}}$-basis for $I\otimes B$ and $V^{*}\otimes W$ is a ${\mathbbm{k}}$-basis for $V\otimes W^{*}$. Also, $V^{*}\otimes W^{*}$ is a ${\mathbbm{k}}$-basis of $R=A\otimes B$. Let $f\in I^{(h)}$ then $f=\sum_{i=1}^{n}c_{i}v_{i}$, where $c_{i}\in{\mathbbm{k}}\setminus\{0\}$ and $v_{i}\in V$. Now
@@
The last equality follows from [HNTT20, Lemma 3.1]. Let $f\in I^{(h)}\subseteq A$ be an arbitrary element, and consider $f\otimes_{k}1\in(I+J)^{(h)}\subseteq(I+J)^{r}=\sum_{i=0}^{r}(IR)^{i}(JR)^{r-i}$. Since $1\notin W$ then $v\otimes 1\notin A\otimes J$. Thus none of the $c_{i}v_{i}\otimes 1$ appearing in the expression of $f\otimes 1$ will be in $(IR)^{j}\cap(JR)^{r-j}$, where $11\leqslant j\leqslant r-$ 1. Hence $f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(IR)^{h}$. Now $(IR)^{h}\cong I^{h}\otimes B$ and thus $f\otimes 1\in(IR)^{h}$ implies that $f\in I^{h}$. This completes the proof. ∎
The following generalizes [BHJT21, Lemma 2.5] to finitely generated algebra domains. The proof of [BHJT21, Lemma 2.5] depends on the binomial summation formula and numerical inequalities, which holds in general domains as well, so we omit the proof.
@@
The proof has the same spirit as [BHJT21, Theorem 2.6] and it follows from Lemma 4.1, Lemma 4.3 and Lemma 4.4. ∎
Now we give the result regarding the summation formula for resurgence of ideals in finitely generated ideals in ${\mathbbm{k}}$ algebras, where ${\mathbbm{k}}$ is algebraically closed field.
@@
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field ${\mathbbm{k}}$ which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. Set $I+J=IR+JR$, where $R=A\otimes B$. Then
@@
We give a sketch of the proof as it follows the same pattern as [KNT24, Theorem 2.3]. First note that by Lemma 4.1, $\rho(IR)=\rho(I)$ and $\rho(JR)=\rho(J)$. The binomial summation formula for symbolic powers hold for ideals in noetherian ${\mathbbm{k}}$-algebras (see [HJKN23, Theorem 4.1]. The other numerical inequalities in [KNT24, Theorem 2.3] follows straightforward. Hence the required result follows. ∎

### v3

@@ 2. Symbolic Powers and Resurgence via flat extension
Now by Lemma 2.4, we have
$\displaystyle\bigcup_{\operatorname{Ass}{\mathfrak{q}}\in\left(I^{(n)}B\right)}{\mathfrak{q}}\subset\bigcup_{{\mathfrak{q}}\in\operatorname{Ass}(IB)}{\mathfrak{q}}=\bigcup_{{\mathfrak{p}}\in\operatorname{Ass}(I)}{\mathfrak{p}}B.$
Therefore whenever $s\in W$ then $s\not\in{\mathfrak{q}}$, where $\mathfrak{q}$} is an associated prime of $I^{(n)}B$. By [atiyah1994introduction, Proposition 4.4(iii)] we get
@@
Note that $I^{(m)}\subset I^{(m)}B\cap A\subset IB^{r}\cap A=I^{r}B\cap A=I^{r}$, the last equality holds since $I^{r}$ is a contraction of some ideal. ∎
@@
Note that if placing $\operatorname{Ass}^{*}(IB)=\{{\mathfrak{p}}B~|~\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}^{*}(I)\}$ by $\operatorname{Ass}(IB^{N})=\{{\mathfrak{p}}B~|~\text{ for all }{\mathfrak{p}}\in\operatorname{Ass}(I^{N})\}$, then the theorem 2.5 is only true up for $1\leqslant n\leqslant N$.
If one uses the minimal prime definition of the symbolic powers, then it is enough to assume that $\operatorname{Min}(IB)=\{{\mathfrak{p}}B~|~{\mathfrak{p}}\in\operatorname{Min}(I)\}$.
The following example shows that the condition in Theorem 2.5 that the associated primes are contraction of the ideals is necessary.
Example 2.7.
Consider the natural map $\phi:{\mathbb{Z}}[x]\to{\mathbb{Q}}[x]$. Note that it is a flat map. Take $I=\langle x,2\rangle\subset{\mathbb{Z}}[x]$. Then $I$ is not a contraction of any ideals of ${\mathbb{Q}}[x]$. Since $I$ is a maximal ideal, then $I^{(n)}=I^{n}$ for all $n\geqslant 1$. But $I{\mathbb{Q}}[x]=\langle x\rangle$. Hence $(I{\mathbb{Q}}[x])^{(n)}=\langle x^{n}\rangle$ for all $n\geqslant 1$. But $I^{(n)}{\mathbb{Q}}[x]=\langle x\rangle$ for all $n\geqslant 1$. Thus $(I{\mathbb{Q}}[x])^{(n)}\neq I^{(n)}{\mathbb{Q}}[x]$ for all $n\geqslant 2$.
@@
In Theorem 2.5 we extend [GHMNmatroid, Theorem 3.6], [Akessesh, Theorem 4], and [walker2018uniform, proposition 2.1] by replacing the faithfully flat condition with flat condition. Note that assumptions in Theorem 2.5 implies all these theorems and proposition, but not vice versa. We now give simple examples satisfying assumptions in Theorem 2.5, but violating the faithfully flat condition required for all theorems and proposition we just mentioned.
Example 2.8.
Consider the map $\phi:{\mathbb{Z}}\to{\mathbb{Z}}\left[1/\penalty 50p\right]$, where $p$ is a prime number, then this map is obviously flat, since $I\otimes_{{\mathbb{Z}}}{\mathbb{Z}}\left[1/\penalty 50p\right]\to{\mathbb{Z}}\otimes_{{\mathbb{Z}}}{\mathbb{Z}}\left[1/\penalty 50p\right]$ is always injective for any ideal $I\subset{\mathbb{Z}}$. However, it is not faithfully flat since $Spec({\mathbb{Z}}\left[1/\penalty 50p\right])\to Spec({\mathbb{Z}})$ is not surjective for closed points, in particular, $(p)$ is the closed point. Finding any ideal $I$ in $\mathbb{Z}$} such that $(p)$ is not an associated prime of $I$, then $I^{(n)}{\mathbb{Z}}\left[1/\penalty 50p\right]=(I{\mathbb{Z}}\left[1/\penalty 50p\right])^{(n)}$. This example is also a good comparison to Example 2.7, since any primes in $\mathbb{Z}$} is a contraction of the same primes in ${\mathbb{Z}}\left[1/\penalty 50p\right]$ except $(p)$. This example can be extended to PID and its prime ideal, since prime ideal (other than zero ideal) is also a maximal ideal, then it can be shown that map from PID to itself with reversing a prime ideal is flat but not faithfully flat by the same reason.
Example 2.9.
@@
Let $A$ be a ring having at least two different prime ideals such that one is not contained in another one, and $B=A_{\mathfrak{p}}$ be localization of $A$ at an arbitrary prime $\mathfrak{p}$. It is a well-known fact that localization is an exact functor, so $B$ is a flat $A$-module, and prime ideal $\mathfrak{p}$ in $A$ is the contraction of $\mathfrak{p}B$. Note that $\mathfrak{p}B$ is the only prime in $B$, then $\operatorname{Spec}(B)\rightarrow\operatorname{Spec}(A)$ is never surjective, so $B$ is not faithfully flat as a $A$-module according to [atiyah1994introduction, ex 3.16].
@@
One can take $A={\mathbb{Z}}$ and $B$ as $\mathbb{Z}_{(p)}$ where $p$ is a prime number. Take $I=(p^{m})$ for some $m\geqslant 1$. Then $I$ satisfies assumptions in Theorem 2.5. Furthermore, letting $B$ be any localization of $A$ will meet the assumptions in Theorem 2.5. This tells us that the stalk of $\operatorname{Spec}A$ gives us enough information of symbolic power of an ideal generated by the corresponding point in $\operatorname{Spec}A$. Another example would be take $A={\mathbb{C}}[t]$ and $B={\mathbb{C}}[t]_{(p(t))}$, where $p(t)$ is an irreducible polynomial in ${\mathbb{C}}[t]$.
Applying our Theorem 2.5, we will have the following corollary, which is crucial for the main theorems in our section 3. We also explain the importance of this corollary in details at the beginning of the section 3.
Corollary 2.10.
@@
Let $\phi:A\to B$ be a flat morphism. If ${\mathfrak{p}}B$ is a prime ideal for all primes $\mathfrak{p}$} in $A$, then $\rho(I)=\rho(IB)$, and $\rho_{a}(I)=\rho_{a}(I)$ for any proper ideal $I$.
@@
The proof follows directly from Theorem 2.5. Note that if ${\mathfrak{p}}B$ is prime for all primes $\mathfrak{p}$} in $A$, then $\phi$ is faithfully flat and $IB\cap A=I$ for all ideals $I$ in $A$. Indeed, by [matsumura1989commutative, theorem 7.2], if for every maximal ideal $\mathfrak{m}$ in $A$ we have $\mathfrak{m}B\neq B$, then $\phi$ is faithfully flat. ∎
Morphisms above with the property that that extend every prime ideal as an prime ideal, also known as morphims with prime extension property, see [hochsterprime]. We will explain the geometric meaning of flat and prime extension, but before that, we need to recall the definition of fiber over points first.
@@
[…]
This proposition implies that if $A$ is reduced, and $\phi$ has prime extension property, then $B$ is also reduced. Indeed, if $A$ contains the zero ideal as a prime ideal, so does $B$. If $\phi$ is a prime extension, then every $B_{\mathfrak{p}}$ is an integral domain. We are ready to give a geometric example if we recall one more lemma.
@@
[stacks-project, Lemma 10.39.5] Let $M$ be an $R$-module, then $M$ be a flat $R$-module if only if for every ideal $I\subset R$ the map $I\otimes_{R}M\rightarrow R\otimes_{R}M$ is injective.
@@
We identify closed points of $A$ with elements of $\mathbbm{k}$}, and obviously the (tc)(t-c) in ${\mathbbm{k}}[t,x,y]/(y-x^{2}-tx)$ is irreducible for every $c\in{\mathbbm{k}}$, so $\phi$ is also a prime extension. The fiber $\operatorname{Spec}B_{(t-c)}$ is given by the plane curve $y=x^{2}+cx$ in $\mathbb{A}_{\mathbbm{k}}^{2}$, which is an irreducible variety, and those plane curves are even homeomorphic to each other.
[…]
@@ 3. Resurgence of sum of ideals
First note that in this section, we assume that all rings and algebras are regular. We summarize current results of resurgence number of sums of ideals for polynomial rings over a field, our goal is to prove these results for sums of finitely generated graded $\mathbbm{k}$}-algebras which are integral domains, where $\mathbbm{k}$} is algebraically closed. In [fiberprojective, Theorem 2.6], it has been proved that if $I$ and $J$ are ideals in different polynomial rings over a same field, then
@@
In [sharpbound, Theorem 2.3], it has been proved that
@@
In our paper, [fiberprojective, Theorem 2.6] is generalized to Theorem 3.5, and [sharpbound, Theorem 2.3] is generalized to Theorem 3.6. The generalization is not trivial. First, it is obvious that the maps from two polynomials rings $A$ and $B$ over a field $\mathbbm{k}$} to $A\otimes_{{\mathbbm{k}}}B$ preserve resurgence numbers. However, for our generalization, we need to apply Corollary 2.10 to achieve that resurgence number of an ideal is same as resurgence number of its extension. Second, in the proof of Lemma 3.3, we use vector space method, while Gröbner basis method was used in the polynomial case but could not be used in $\mathbbm{k}$}-algebra case.
In Lemma 3.1, we will specify that under proper conditions resurgence of the extension is the same as the resurgence of the ideals.
@@
Let $A$ and $B$ be finitely generated $\mathbbm{k}$} algebras, where $\mathbbm{k}$} is algebraically closed, which are integral domains. If $\phi:A\to R$ is a canonical ring homomorphism, where $R=A\otimes_{{\mathbbm{k}}}B$, then $\phi$ is flat and has prime extension property. Furthermore, $\rho(IR)=\rho(I)$ and $\rho_{a}(IR)=\rho_{a}(I)$ for any proper ideal $I$ from $A$ (or $B$).
@@
Now $\left(A/\mathfrak{p}\right)\otimes_{k}B$ is an integral domain by [algvar, Lemma 1.54]. It is a well-known fact that an ideal is prime if only if the quotient ring is an integral domain, so $\mathfrak{p}R$ is prime, and thus, $\phi$ has prime extension property. The last equality follows from Corollary 2.10. ∎
The next lemma also extends [fiberprojective, Lemma 2.4]. The proof of [fiberprojective, Lemma 2.4] depends on divisibility in the polynomial ring. While working with general domains, we treated them as vector spaces and used properties of basis. Let us recall the following definition from [guardo2013asymptotic].
Definition 3.2.
Let $I$ be an ideal in a regular ring. Then $\lim_{t\to\infty}\sup\rho(I,t)$ is defined as follows:
$\rho(I,t)=\sup\left\{\dfrac{m}{r}~|~I^{(m)}\not\subseteq I^{r},m\geqslant t,r\geqslant t\right\},\text{ and }\rho^{\lim\sup}_{a}(I)=\lim_{t\to\infty}\sup\rho(I,t).$
@@
Take $\theta\in{\mathbb{Q}}$ such that $\theta>\rho_{a}^{\lim\sup}(I+J)$. Then there is a $s_{0}\in{\mathbb{N}}$ such that for all $t\geqslant s_{0},\theta>\rho(I+J,t)$, where $\rho(I+J,t)=\sup\left\{\dfrac{m}{r}~|~I^{(m)}\not\subset I^{r},m\geqslant t,r\geqslant t\right\}$. Thus for any $h,r\in{\mathbb{N}}$ so that $h,r\geqslant t\geqslant s_{0}$ and $\dfrac{h}{r}\geqslant\theta$, one has $(I+J)^{r}\supset(I+J)^{(h)}\supset I^{(h)}$. Up to this point the proof idea is same as [fiberprojective, Lemma 2.4]. Now we need to show that $I^{(h)}\subseteq I^{r}$.
Let $V$ be a $\mathbbm{k}$} basis of $I$, which extends to $V^{*}$ a $\mathbbm{k}$}-basis of $A$ containing $1$. Let $W$ be a $\mathbbm{k}$}-basis of $J$, which extends to $W^{*}$, a $\mathbbm{k}$}-basis of $B$ containing $1$. Then $V\otimes_{{\mathbbm{k}}}W^{*}$ is a $\mathbbm{k}$}-basis for $I\otimes_{{\mathbbm{k}}}B$ and $V^{*}\otimes_{{\mathbbm{k}}}W$ is a $\mathbbm{k}$}-basis for $V\otimes_{{\mathbbm{k}}}W^{*}$. Also, $V^{*}\otimes_{{\mathbbm{k}}}W^{*}$ is a $\mathbbm{k}$}-basis of $R=A\otimes_{{\mathbbm{k}}}B$. Let $f\in I^{(h)}$ then $f=\sum_{i=1}^{n}c_{i}v_{i}$, where $c_{i}\in{\mathbbm{k}}\setminus\{0\}$ and $v_{i}\in V$. Now
@@
The last equality follows from [ha2020symbolic, Lemma 3.1]. Let $f\in I^{(h)}\subseteq A$ be an arbitrary element, and consider $f\otimes 1\in(I+J)^{(h)}\subseteq(I+J)^{r}=\sum_{i=0}^{r}(IR)^{i}(JR)^{r-i}$. Since $1\notin W$ then $v\otimes 1\notin A\otimes J$. Thus none of the $c_{i}v_{i}\otimes 1$ appearing in the expression of $f\otimes 1$ will be in $(IR)^{j}\cap(JR)^{r-j}$, where $1\leqslant j\leqslant r-1$. Hence $f\otimes 1=\sum_{i=1}^{n}cv_{i}\otimes 1\in(IR)^{h}$. Now $(IR)^{h}\cong I^{h}\otimes B$ and thus $f\otimes 1\in(IR)^{h}$ implies that $f\in I^{h}$. This completes the proof. ∎
The following generalizes [fiberprojective, Lemma 2.5] to finitely generated algebra domains. The proof of [fiberprojective, Lemma 2.5] depends on the binomial summation formula ([bionomial, Theorem 3.4]) and numerical inequalities, which holds in Noetherian algebras over a field as well, so we omit the proof.
@@
The proof has the same spirit as [fiberprojective, Theorem 2.6] and it follows from Lemma 3.1, Lemma 3.3 and Lemma 3.4. ∎
@@
Let $A$ and $B$ be finitely generated algebras over an algebraically closed field $\mathbbm{k}$} which are integral domains, and let $I\subset A$ and $J\subset B$ be nonzero proper homogeneous ideals. Set $I+J=IR+JR$, where $R=A\otimes_{{\mathbbm{k}}}B$. Then
@@
We give a sketch of the proof as it follows the same pattern as [sharpbound, Theorem 2.3]. First note that by Lemma 3.1, $\rho(IR)=\rho(I)$ and $\rho(JR)=\rho(J)$. The binomial summation formula for symbolic powers hold for ideals in noetherian $\mathbbm{k}$}-algebras (see [ha2023binomial, Theorem 4.1]. The other numerical inequalities in [sharpbound, Theorem 2.3] follows straightforward. Hence the required result follows. ∎
Remark 3.7.
To see why we need these conditions on $A$ and $B$ from Theorem 3.5, first note that, if $A$ (as a field extension over $\mathbbm{k}$}) is not finitely generated as a $\mathbbm{k}$}-algebra, then $A\otimes_{\mathbbm{k}}A$ may not a Noetherian ring by [vmos, Theorem 11]. For an example, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is not a Notherian ring. Therefore, the condition being finitely generated is necessary. We also need $A$ and $B$ to be integral domains as $\mathbbm{k}$}-algebras where $\mathbbm{k}$} is algebraically closed because of [algvar, Lemma 1.54], this will make $A\otimes_{\mathbbm{k}}B$ to be an integral domain. A famous example for violating the algebraically closed condition, then we do not have integral domain, is $\mathbb{C}\otimes_{\mathbb{R}}\mathbb{C}$. Moreover, note by [Regulartensor, Theorem 6(e)], $A\otimes_{{\mathbbm{k}}}B$ is regular if $A$ and $B$ are regular.